IBM’s current stock price is $134, and its last dividend was $6.70. In view of IBM’s strong financial position and its consequent low risk, its required rate of return is only 11 percent. If dividends are expected to grow at a constant rate, g, in the future, and if rs is expected to remain at 11 percent, what is IBM’s expected stock price five years from now?
Answer #1
Current Dividend $         6.70
Rate of return 11.00%
Growth Rate g
Share price $     134.00
With the given information, first, we should calculate the growth rate and then we can proceed to assess
the position 5 years after.
Share Price =Current Dividend*(1+Growth rate)/(Rate of return-Growth Rate)
134 =6.7*(1+g)/(0.11-g)
(1+g)/(0.11-g)= 134/6.7
(1+g)/(0.11-g)= $       20.00
(1+g)= 20*(0.11-g)
(1+g)= 20*0.11- 20*g)
(1+g)= 20*0.11- 20*g)
1+g= 2.2 - 20*g
g+20 g= 2.2-1
21*g= 1.2
g= 1.2/21
g= 5.71%
Dividend after 5 years 6.70*(1+5.71%)^5
Dividend after 5 years $         8.85
Share price after 5 years= =Current Dividend*(1+Growth rate)/(Rate of return-Growth Rate)
Share price after 5 years= =8.85*(1+5.71%)/(11%-5.71%)
Share price after 5 years= $     177.00